Summary of the 3mm GMVA Session of May 17-22, 2012 -------------------------------------------------- Stations participating: EB, On, PV, PdB, Yb, VLBA Metsahovi did not participate 3 KVN telescopes participated for part of the time On and Yb have LCP, other have LCP+RCP The ftp fringe test on May 17 showed fringes to all European stations. Pico Veleta recorded part of the time with MK5A (analog) and MK5C (digital DBBC) Effelsberg recorded part of the time with MK5A (analog) and MK5C (digital DBBC) The overall weather conditions in Europe were pure. This is not unexpected, since this session was schedule to late in the year and outside the optimum observing time window. The use of large 8packs causes problems, at Pico two 6 TByte diskpacks failed when they reached a filling of 50 percent. The same may have appeared at at least one VLBA station (PT). Effelsberg: ---------- Receiver temperature appears to be 160-170 K higher than usual (130 K). Weather was variable, a mixture of clear sky periods, followed by clouds and thunderstorm, then clear gain. May 17: thin cloud layer, observing ok May 18: ok until 12h00, then heavy clouds, wind May 19: overcast, later thin cirrus, patchy blue sky May 20: overcast, later scattered clouds, blue sky May 21: very cloudy May 22: thin layer, clear sky Pico Veleta: ------------ 0-baseline test between analog equipment and DBBC at Pico Veleta was successful (H. Rottmann). At start: The sky is not very clear, the Tsys is about 140K. May 18: 1h30-5h30 UT lost, high wind May 19, morning: partial cloudy, wind high May 18, 22h30 - May 19, 3h45 lost, high wind May 19, 10h00 - 11h15, both MK5 recorder hang, exchange disk pack May 19, 20h00- May 21, 08h45 lost, bad weather, snow+wind May 20, 12h00: H-maser out of lock since night of May 19. May 21, 09h00 until end ok PdBure: -------- At start weather is overcast and degrading. Start with 4 antennas, later get 5. May 17, morning: Plateau de Bure observing with 5 antennas under light clouds and wind close to our limits. May 18, 2h30-10h30 snow, not observing from 10h45 onwards ok May 19, most of it ok until the end of c121a May 20, morning: started c121b on time under cloudy conditions that improved later. Since May 20 5:10 UT stopped due to high winds, snow and rain. No further observing until end of c121b on May 22, 16h30. Yebes: ------- May 17: first Scans (1-6) badm after that ok, weather good May 18: some high wind May 19: 04-07h20 UT lost, wrong date and time of field system May 20 to end: cloudy weather, showers, some wind Onsala: ------- started in rain, later better, but during the whole session the weather wasn't very good. Lock at the Tsys plots on vlbeer for detailed info. day 138: 17h-24h UT high Tsys day 139: 0-24h Tsys ok day 140: 0-12h Tsys ok, 12-24h high Tsys day 141: 0-10h Tsys ok, 10-18h higb Tsys day 142: Tsys ok Here some weather info from the logs: c112a: 2012.139.01:02:10.54;"cc=100%, ws=13,3m/s, T=7,6C 2012.139.15:01:27.43;"cc=0%, ws=8,6m/s, T=9,5C 2012.139.18:04:14.18;"cc=0%, ws=6,2m/s, T=9,9C 2012.139.19:02:54.32;"cc=10%, ws=4,7m/s, T=10,1C 2012.139.20:04:20.91;"cc=30%, ws=3,4m/s, T=10,3C 2012.139.20:58:50.69;"cc=30%, ws=4,0m/s, T=10,0C 2012.139.22:02:25.01;"cc=30%, ws=4,4m/s, T=10,2C 2012.139.23:24:16.40;"no rain, ws=4.1m/s, T=9.0C 2012.140.00:09:10.45;"no rain, ws=3.6m/s, T=9.2C 2012.140.00:58:22.95;"no rain, ws=4,4m/s, T=9.5C 2012.140.01:51:03.39;"cc=100%, ws=5,2m/s, T=9.7C 2012.140.03:14:42.49;"cc=90%, ws=4,5m/s, T=9.3C 2012.140.03:43:14.58;"cc=90%, ws=4,8m/s, T=9.3C 2012.140.06:08:07.11;"cc=80%, ws=4,7m/s, T=12.2C 2012.140.07:03:00.91;"cc=70%, ws=5,3m/s, T=12.2C 2012.140.08:08:20.43;"cc=70%, ws=6,4m/s, T=14.1C 2012.140.09:06:28.47;"cc=70%, ws=8,7m/s, T=15.3C 2012.140.10:00:56.27;"cc=90%, ws=4,3m/s, T=12.1C 2012.140.10:59:42.86;"cc=100%, ws=6.3m/s, T=12.4C 2012.140.13:06:50.75;"cc=90%, ws=4,7m/s, T=14.3C 2012.140.14:15:21.76;"cc=100%, ws=4,7m/s, T=14.8C 2012.140.15:10:16.30;"cc=100%, ws=5,6m/s, T=15.4C 2012.140.16:00:31.46;"cc=100%, ws=4.1m/s, T=14.6C 2012.140.16:58:08.54;"cc=100%, ws=5.7m/s, T=13.3C 2012.140.17:01:58.69;"RFI on if21 & if2 2012.140.18:08:17.37;"cc=100%, ws=5.7m/s, T=12.9C 2012.140.18:58:15.34;"cc=100%, ws=4.7m/s, T=13.3C c112b: 2012.142.22:06:35.32;"cc=10%, ws=6.9m/s, T=19.2C 2012.142.23:01:24.04;"cc=50%, ws=7.0m/s, T=18.7C 2012.143.00:02:59.59;"cc=50%, ws=5.7m/s, T=17.8C 2012.143.00:57:38.99;"cc=50%, ws=6.5m/s, T=17.3C 2012.143.01:55:58.78;"cc=50%, ws=6.4m/s, T=16.9C 2012.143.03:03:08.76;"cc=50%, ws=6.3m/s, T=16.2C 2012.143.03:59:51.69;"cc=30%, ws=6.1m/s, T=16.0C 2012.143.05:00:07.11;"cc=10%, ws=6.8m/s, T=16.7C 2012.143.07:03:19.93;"cc=10% 2012.143.08:57:07.46;"cc=10% 2012.143.14:12:08.89;"cc=90%, ws=5,2m/s, T=26,4C 2012.143.14:59:40.82;"cc=100%, ws=5,5m/s, T=26,5C 2012.143.15:55:50.51;"cc=100%, ws=5,5m/s, T=25,7C KVN: ---- 3 stations participating to record about 1 Tbyte of data in selected scans for VLBI fringe test: Ku=Ulsan, YS=Yonsei, TN=Tamna No major problems known, all stations apparently working properly. VLBA: ----- VLBA: Observations were interrupted by two schedule breaks (a 4 hrs ToO on SgrA* on May 17 (8-12h30 UT) and the HSA experiment BM352 on May 19 1230-2130 UT. By human error PT erased in experiment c121b one 8 pack. All previous recorded data on that 8pack are likely to be lost. Detailed logs are given below: Project: C121A Observer: Krichbaum, T. Project type: VLBI Obs filename: c121acrd.* Date/Day: 2012MAY17/138 Ants Scheduled: NL FD LA PT KP OV BR MK =UT-Time===Comment===============================================MF#===%AD==AMD= Operator is Paul Padilla UT1-UTC: -0.56, VClock: 0.023, GPS delta: -7.8 %NL and LA 3mm receivers running a bit warm. CRYO Tsys values will be elevated. 2326 Begin at all sites except MK 2326-1127 *OV 7mm LCP is weak (C131067) FE 25 540 2343-0143 *PT,LA not running script in the VLBA Monitor SW 200 240 screen. Will attempt to restart executor to see if that will get it going. 0000 Date is 2012MAY18/139. 0007 *Ok, They're still not showing up on the GUI, although the execInfo command indicates they are running the jobs. On the other hand, the Tsys at PT still don't look right. 0011-0016 *NL stopped sending back monitor data, rebooting VME SW 100 5 0014 *NL came back up out of point mode 0045 Got in contact with Jim. He will be by in a little while to help with the PT/LA problem 0047 *Looking closer at the execInfo, it appears that the script is showing up in the "history" section for both sites instead of the "active" section. 0112 Jim came by, but was unfortunately no help. Still waiting on word from Walter or Matthias. 0142 *PT finally showed up after Walter played around with it for a little while. 0143 *LA also showed up, now. 0143-1127 *LA 3mm RCP is dead (C130439). FE 25 506 0338-1230 *FD communications is going in and out. Rebooting HWC 100 532 the VME to see if that is the issue (C131278). 0346 *FD still not happy. Trying to reboot Mark5, now. 0415 MK Begins 0420 *FD still having issues, calling the Site Tech to see if he can see it better on his end closer to the site. 0430 &FD tech heading to site to investigate C131278 0545 New operator: T. Perreault 0551 *FD had the site tech park the antenna and reset the dial backup modem (which did not connect). Antenna is in a safe condition. Emailed network support in Charlottesville, should see the email around 1200 - 1400 UT. 0614-0620 *MK program is not showing up in the VLBA monitor SW 100 6 GUI; logged into the control computer and attempted to run execInfo, no response. Did a RestartExecutor command, and tried the execInfo again, no response. Did a reboot of the control computer. 0620 *MK was able to get program started on the control computer following reboot. 0730-1230 Suspending C121A for TOO observation BE061C. SCH 800 2400 1230 Resuming C121A 1230-0116 *FD communication is still out with the site. HWC 100 766 (C131278) 1339 *Networking support reports that problem at FD seems to be with the station computer -- can ping other devices such as the MARK5A recorder on the local net. Network support is talking with the site tech. 1352 PT wea: cloudy, no precipitation 1358 LA wea: scattered high clouds, no precipitation 1417 Operator is A. Sowinski. 1440 BR wea: high thin clouds, dry. 1457 OV wea: clear, dry. 1515-1516 *PT HC1 formatter errors. C121522. FMTR 100 1 1531 KP wea: scattered high clouds, dry. 1850-1851 *KP Autostow, repointed. COM 100 1 1924-1934 *LA Site tech working on antenna. FE 100 10 1925-1957 *NL Antenna stowed. VME hung and rebooted SW 100 32 by site tech. Due to internet problem, local control of antenna was both necessary and difficult. 2147 New Operator: Paul Padilla 0000 Date is 2012MAY19/140. 0107-0122 *OV not sending back monitor data, rebooting VME. SW 100 15 0118 *OV came back in STANDBY and with Mark5 OFFLINE 0126-0133 *LA not recording - there is a slew of "bank read REC 100 7 error"s for the Mark5. Rebooting it. 0530 New operator: T. Perreault 0600-1127 *OV 2-16 GHz synthesizers 2 and 3 are not locking PS 100 327 (also synthesizer 1, which is not used at 7mm & 3mm) at commanded frequency. Suspect that a power supply has died since it is the only thing common with all three synthesizers. Checked rack power supplies, the A-rack +15V PS 1 and 2 are both low, but the synthesizers are in the B-rack. Possible that the power supplies are affecting the synthesizers. (C131280) 1057 MK observing ends 1127 End. Downtime Summary: Total downtime : 5388 min Percentage downtime of observing: 31.2% Average downtime per hour : 18.7 min VERY LONG BASELINE ARRAY OBSERVING LOG -------------------------------------- Project: C121B Observer: Krichbaum, T. Project type: OTHER Obs filename: c121bcrd.* Date/Day: 2012MAY19/140 Ants Scheduled: NL FD LA PT KP OV BR MK =UT-Time===Comment===============================================MF#===%AD==AMD= Operator is A. Sowinski. UT1-UTC -0.56; gps delta and vclock okay. 2226 Begin. 2226-1627 *LA 3mm RCP dead. C130439. FE 50 1980 2226-1627 %BR 7mm LCP Tsys higher than normal. C126577. 2226-1627 %NL 3mm receiver slight warm (37K). CRYO 0000 Date is 2012MAY20/141. 0045-0157 %OV Flaky power supply causing Syn1 and Syn3 PS 50 36 problems. C131280. 0147 New operator is Betty Ragan. 0157-0216 *OV out of point. Site tech is working on the PS 100 19 power supply. 0358-0413 *OV 15V A-rack power supply was doing better PS 75 11 after the site tech cleaned out the connectors and re-seated it, but has been steadily losing power, and now synthesizers 2 and 3 are again not locking. (C131280) 0413-1511 *OV syn 1 is now not locking, either. Removing PS 100 2098 it from the observation. (C131280) 1345 Operator is A. Sowinski. 2238-2247 *NL VME partial communications loss. Rebooted. HWC 100 9 0000 Date is 2012MAY21/142. 0144 New operator is Betty Ragan. 1302-1304 *KP autostowed. Power glitch. COM 100 2 1338 Operator is A. Sowinski. 1343 LA wea: overcast, low clouds, dry. 1400 PT wea: clear, light wind, dry. 1406 BR wea: overcast high clouds, dry. 1408-1409 *BR HC1 formatter errors. FMTR 100 1 1507 OV Site tech is testing a different power supply. Putting antenna back into project. 1524 KP wea: clear, dry. 1842 %BR Raining. 1930 MK wea: clear, easterly winds. 1934-1936 *NL Site tech swapping D-rack power supply. 1941 %FD Raining. 0730-2041 *PT Mark5 module MPI+1210 error: Problem decoding REC 100 791 module, try reinserting. Module showed zero usage, so I thought the site tech had just inserted it. I erased it to try to clear the error, which worked. However the module had actually been recording since yesterday, so a lot of data was lost. I set this time range to include all of it. 2145 New Operator is Alan Kerr 2146-1627 *OV out. Site Tech checking power supply. C131280 PS 100 1121 UPDATE: ps voltage is low. cannot keep any synthesizers at proper levels. removing from observation. 2357-0011 *KP lost communications. vme rebboted. SW 100 14 0000 Date is 2012MAY22/143. 0004-0006 *PT out of point. power glitch. COM 100 2 0537 New Operator: Paul Padilla 1326 New operator: T. Perreault 1345 PT wea: clear, no precipitation 1401 LA wea: clear, no precipitation 1431 BR wea: broken high clouds, no precipitation 1455 OV wea: clear, no precipitation 1528 KP wea: clear, no precipitation 1627 End. Downtime Summary: Total downtime : 6084 min Percentage downtime of observing: 19.2% Average downtime per hour : 11.5 min